The volume of a gas is 36.0 ml at 10.0 degree C an
Answer #1

35)
Vi = 36.0 mL
Ti = 10.0 oC = (273 + 10.0) K = 283 K
Pi = 4.50 atm

Vf = 85.0 mL
Tf =?
Pf = 3.50 atm

use:
Pi*Vi/Ti = Pf*Vf/Tf
4.50* 36.0 / 283 = 3.50*85.0/Tf
Tf = 519.7 K
= (519.7 - 273) oC
= 246.7 oC

Answer: e

36)
V = 125 mL = 0.125 L
P = 75.0 atm
T = 25 oC = (25 + 273) K = 298 K

use:
P*V = n*R*T
75.0 * 0.125 = n*0.0821*298
n = 0.383 mol
Answer: d

I am allowed to answer only 1 question at a time